One pipe fills a storage tank in 30 hours. A second pipe fills the same tank in 45 hours. When a third pipe is added and all three pipes are used to fill the tank it only 9 hours. Find how long it takes the third pipe to do the job.
Answer by lwsshak3(11628)

One pipe fills a storage tank in 30 hours. A second pipe fills the same tank in 45 hours. When a third pipe is added and all three pipes are used to fill the tank it only 9 hours. Find how long it takes the third pipe to do the job.
**
Work rates:
1st pipe: 1/30
2nd pipe: 1/45
3rd pipe: 1/x
All three pipes used: 1/9
..
sum of individual rates = rate when all working together
..
1/30+1/45/+1/x=1/9
1/x=1/9-1/30-1/45
LCD:90x
90=10x-3x-2x=5x
5x=90
x=18
ans:
It would take the 3d pipe 18 hours to do the job alone